# PixFoundation: Are We Heading in the Right Direction with Pixel-level Vision Foundation Models?

### v1

@@ 4.1 Experimental Setup
Evaluation benchmarks, protocols and metrics. PixMMVP is composed of 300 images paired with questions, choices, referring expressions and segmentation masks, while PixCV-Bench has 1,438 images with their corresponding annotations similarly. On each benchmark we evaluate the VQA and visual grounding capabilities following three probing techniques and reporting their metrics. The first probing is to evaluate the VQA ability, where the accuracy is computed using GPT-4o following (Tong et al., 2024b) as, . If the model generates a segmentation without explicitly asking it to, it is evaluated with respect to the ground-truth referring segmentation in terms of mean intersection over union as, . The second probing prompts the model to identify the referred expression then evaluates the mean intersection over union reported as, . The third probing following (Tong et al., 2024a) instructs the model to generate a single option letter and evaluate the accuracy directly without GPT-4o, reported as, . There is a need for the first probing since some of the recent pixel-level MLLMs face challenges in following instructions. We evaluate the score of each model, , which is the harmonic mean across the maximum of both pixel-level visual grounding and VQA,
We mainly focus on evaluating four state-of-the-art pixel-level MLLMs; LISA (Lai et al., 2024), GLAMM (Rasheed et al., 2024), OMG-Llava (Zhang et al., 2024b) and Llava-G (Zhang et al., 2024a). For GLAMM we use two variants; the original model (GLAMM) and the one fine-tuned for region-level captioning, (GLAMM-RegCap). We use Hugging Face for these models weights, refer to Appendix A.
Baselines and upper bound implementation details. We evaluate: (i) the attend and segment (a+s), (ii) the oracle selection relying on the highest intersection over union in selecting the correctly predicted masks (PixFoundation), and (iii) the automatic selection, (PixFoundation). These are implemented on top of three base MLLMs which are, Llava 1.5 (7B, 13B) (Liu et al., 2024) and Cambrian-1(8B) (Tong et al., 2024a). The automatic selection is implemented using the Cambrian-1 (8B) model, yet it can be further improved using GPT-4o, refer to Appendix A for more details.
@@ 4.2 Are the current pixel-level MLLMs heading in the right direction?
In order to answer this, we evaluate each of these pixel-level MLLMs capability in VQA in challenging tasks. Additionally, we evaluate their ability to visually ground the objects of interest in these questions.
PixMMVP Benchmark. Table 1 shows the results on the challenging PixMMVP benchmark. From the accuracy of VQA, MLLMs that are not trained with pixel-level grounding surpass their pixel-level counterpart with up to 16%. The best in pixel-level MLLMs score in this aspect is GLAMM-RegCap (Zhang et al., 2024b) yet it has degraded ability to generate segmentation. On the other hand, when looking at pixel-level visual grounding we find the best model, GLAMM (Rasheed et al., 2024), has a weak ability in VQA or following instructions. Moreover, it shows LISA and Llava-G are incapable of following the instruction to output the option letter reported in, . Looking at the bottom three rows, the oracle confirms that MLLMs that were never trained with pixel-level grounding have the correct grounding within their learned attention maps, refer to Fig. 5. Additional qualitative analysis is in Appendix B. Looking at the final score, , the oracle variant, PixFoundation (7B), outperforms the corresponding best pixel-level MLLMs, OMG-Llava (7B), by a considerable margin while the automatic outperforms it with up to 4%. While the attend and segment baseline (Cao et al., 2024) lags behind our automatic method with more than 6%.
PixCV-Bench Benchmark. Table 2 shows the results on PixCV-Bench. Persistently all the pixel-level MLLMs still lag behind the MLLMs that were never trained with pixel grounding supervision in VQA. Nonetheless, OMG-Llava strikes the right balance in both VQA and pixel-level grounding with the highest score, , within the 7B models. While, GLAMM, shows the best pixel-level grounding yet it is incapable of VQA on this benchmark. Looking at the bottom three rows for the oracle similar findings to PixMMVP emerge. The attend and segment baseline and the automatic PixFoundation baseline, both lag behind the oracle as expected yet they show competitive performance in pixel-level visual grounding, even better than some of the MLLMs trained with pixel-level supervision. Our automatic PixFoundation outperforms the attend and segment using the Llava variants, and is competitive in Cambrian-1. Cambrian-1 is more challenging as it generates more tokens, thus selecting among higher number of masks, refer to Appendix C.
Summary. In summary, pixel-level grounding supervision degrades MLLMs ability in VQA and sometimes even their generalization in grounding. Failure case analysis is provided in Appendix D. We show that MLLMs trained with pixel-level supervision lag behind vanilla MLLMs using simple mechanisms to extract grounding, and the oracle indicates there is an opportunity to improve this. Moreover, we show that grounding might not coincide with the noun phrase most similar to the referred expression, where our oracle upper bound is surpassing the attend and segment.
@@ 4.3 When does grounding emerge in MLLMs?
When - location. Taking into account the powerful performance of the oracle upper bound, it begs the important question of when grounding emerges. We start by looking at when it emerges in terms of the location within the output text. We analyze the word/phrase location with respect to the full output text in terms of a percentage from its total length, (i.e., 0% means the beginning of the text). Accordingly, Fig. 6(a) shows the location percentages histogram, binned at 10%, for the three base MLLMs reporting the oracle selection and evaluating on PixMMVP benchmark using the second probing. In the Llava 1.5 variants, grounding mostly emerges at the last 40% of the output text, while for Cambrian it is mostly at the last 10%.
When - concept. For the second analysis we look into the concept category that the correct output word/phrase corresponds to. The previous assumption in other works is that grounding emerges in exact noun or noun phrase of the object of interest. Except our analysis confirms that this is not necessarily the case. We take the correct noun/noun phrase where the grounding emerges based on the oracle from all the three variants, then we pass it to GPT-4o to request a grouping of these concepts. Specifically, we prompt it with the following “Can you group these noun phrases into categories of abstract concepts. You are provided with pairs of (noun phrase, its full context).”, where the entire output text is the full context. After manual filtration and merging it resulted into six main groups, which are: (i) Color and appearance, (ii) location and position, (iii) object parts, (iv) context and setting, (v) objects and entities, and (vi) State. We then prompt for each of the noun/noun phrase, GPT-4o, to categorize it within these six categories. The histogram of the occurrences of these concept categories is shown in Fig. 6(b). It clearly conveys that in certain scenarios the correct output when grounding emerges can be describing the position or the location of the object of interest not necessarily the exact ground-truth referring expression. While the majority occurs under “Objects and Entities”, there are still percentage of output coinciding with position, color, context, object part, or even the state. Further qualitative results are provided in the Appendix B.
@@
Summary. In summary, we found that grounding emerges in the last 40% of the output text which might indicate that the reasoning of the model to provide a response impacts when the correct visual grounding can occur. More importantly, we show that grounding in MLLMs can emerge in the noun phrase that corresponds to color, position or other characteristics of the object of interest and not necessarily the exact referring expression.
@@ Appendix A Additional implementation details
Models. We also detail the model checkpoints we use for the four pixel-level MLLMs and their variants, retrieved from HuggingFace (Wolf et al., 2019) in Table 3. These also include the model checkpoints used for the base MLLMs that were not trained with pixel-level visual grounding. It is worth noting, that for GLAMM we use two variants (FullScope and RegCap) since their base model (i.e., FullScope) has low performance in the visual question answering task. As such we use the other variant for GLAMM that was fine-tuned for region-level captioning using RefCOCOg and Visual Genome dataset. Furthermore, we provide details on the oracle selection mechanism, where in the case of multiple masks exhibiting the same IoU we select the last occurring one in the when analysis, since we need to select the location of when it occurs. While in the quantitative and qualitative evaluation, we resort to simply not selecting any mask. These occur in few cases when the ground-truth mask is all background, thus any mask IoU will result in zero.
Additionally, we provide details on the SAM model that is used in the three baselines and upper bounds in our benchmarks, where we use the ViT-H variant. Finally, we provide an illustrative example of our automatic selection mechanism with the corresponding predictions on PixMMVP using Llava 1.5 (7B) in Fig. 9. Our automatic selection goes through an iterative process of prompting the selected MLLM, in our case Cambrian-1, with three images overlayed with predicted segmentation to select the best within each group of three. In the final stage the best images are used to prompt the MLLM to select the final selected mask that best describes the object of interest. In the oracle upper bound whenever the model is to be evaluated in a multiple object scenario we take all the possible pairs of the masks then select the best pair based on the highest intersection over union.
@@
| Llava 1.5 (7B) | 4bit/llava-v1.5-7b |
| Llava 1.5 (13B) | 4bit/llava-v1.5-13b-3GB |
@@
Evaluation. We also provide the details on computing the visual question answering accuracy using GPT-4o in the first protocol (Tong et al., 2024b). We use the following prompt: “Given the following question QUESTION, the correct answer is ANSWER. Does the following answer correctly answers the question, answer: RESPONSE? Respond with a Yes/No”.
@@ B.1 Baselines ablation
Attend and Segment (Cao et al., 2024) (Cambrian-1 (8B))
@@ B.3 When does grounding emerge?
We show additional examples on when grounding emerges in multi-modal large language models, specifically in the Llava 1.5 (7B) variant, using the second probing to prompt the model to segment what is in the referring expression. Figures 12, 14, 13, 15 and 16 show the corresponding predicted masks for the grounding that emerged highlighted in red with the maximum attention point as a black dot. Figure 17 shows the aforementioned five examples with the input prompt, the concept category and the noun phrase corresponding to the best grounding using the oracle selection and the full output text. It clearly shows that the correct output token can correspond to location or color but not necessarily the ground-truth referring expression. While some of the noun phrases and their masks from the point prompting of SAM do correspond to what the noun phrase is describing. It is not always the case, example in Fig. 14 “the flame” was not able to highlight the correct object yet it appeared in the noun phrase corresponding to the location “the top”. It also shows that the majority of correct grounding occurs near the end of the output text after the model reasoned for an answer to the question.
Additionally, this type of mining for attention maps can show potential mistakes that MLLMs fall into while performing the grounding to reason on what is in the image. An example on that is in Fig. 15, where the model potentially mistakes a crack in the wood for the minute hand. Figure 16 shows another clear example, where the MLLM thinks it is seeing a tongue highlighting its potential place near the mouth. Yet the image does not really have the tongue out and consequently the error in the grounding lead to an error in the visual question answering. Thus, our PixFoundation serves as an interesting tool to interpret and understand how MLLMs work and reason to produce the final output with the oracle selection as an upper bound. Yet, there is an open question on how to identify the real mistake from the MLLM vs. simple misalignment between the noun phrase and the grounding emerging, since grounding could emerge with phrases describing it or parts of it instead of the object’s noun phrase itself. We leave this for future work.
@@
The minion
the image
its tongue
a blue overalls
a yellow shirt
| Image | Prompt | Concept Category | Noun Phrase | Output |
| 292 | Identify the minion in the image smiling with its tongue out in the scene. | Color & Appearance | a yellow shirt | The minion in the image is smiling with its tongue out, wearing a blue overalls and a yellow shirt. |
[5pt]
1
[5pt]
3
[5pt]
6
[5pt]
161
[5pt]
292
@@ B.5 PixCV-Bench benchmark
Appendix C Analysis on the output length
In this section, we provide additional analysis on the output length on average through PixMMVP dataset using the first and second probing schemes. Specifically, we report the output length as the number of characters in the output, and the number of noun phrases extracted from it. The reason to study this, since it has relation to the number of noun phrases and consequently the number of masks our baselines are selecting among. Table 4 shows the average output length computed across PixMMVP dataset, comparing the three base MLLMs. We notice that Cambrian-1 (8B) generates longer outputs with a considerable margin than Llava variants. Hence, we believe the superiority of the oracle upper bound with Cambrian-1 in the grounding has strong correlation to producing longer outputs with more attention maps to mine and select from, than Llava variants. Nonetheless, it makes it more challenging for the automatic baseline which explains its results on PixCV-Bench with respect to the attend and segment. We believe the use of GPT-4o for the automatic selection instead of Cambrian-1 (8B) should improve these results further.
@@ Appendix C Analysis on the output length
Appendix D Failure Cases Analysis
D.1 Failures in Visual Question Answering
D.2 Failures in Pixel-level Visual Grounding

### v2

@@ 4.1 Experimental Setup
Evaluation benchmarks, protocols and metrics. PixMMVP is composed of 300 images paired with questions, choices, referring expressions and segmentation masks, while PixCV-Bench has 1,438 images with their corresponding annotations similarly. On each benchmark we evaluate the VQA and visual grounding capabilities following three probing techniques and reporting their metrics. The first probing is to evaluate the VQA ability, where the accuracy is computed using GPT-4o following [19] as, . If the model generates a segmentation without explicitly asking it to, it is evaluated with respect to the ground-truth referring segmentation in terms of mean intersection over union as, . The second probing prompts the model to identify the referred expression then evaluates the mean intersection over union reported as, . The third probing following [18] instructs the model to generate a single option letter and evaluate the accuracy directly without GPT-4o, reported as, . There is a need for the first probing since some of the recent pixel-level MLLMs face challenges in following instructions. We evaluate the score of each model, , which is the harmonic mean across the maximum of both pixel-level visual grounding and VQA,
We mainly focus on evaluating four state-of-the-art pixel-level MLLMs; LISA [9], GLAMM [16], OMG-Llava [27] and Llava-G [26]. For GLAMM we use two variants; the original model (GLAMM) and the one fine-tuned for region-level captioning, (GLAMM-RegCap). For details on the models weights, refer to App. A.
Baselines and upper bound implementation details. We evaluate: (i) the attend and segment (a+s), (ii) the oracle selection relying on the highest intersection over union in selecting the correctly predicted masks (PixFoundation), and (iii) the automatic selection, (PixFoundation). These are implemented on top of three base MLLMs which are, Llava 1.5 (7B, 13B) [12] and Cambrian-1(8B) [18]. The automatic selection is implemented using GPT-4o, refer to App. A for more details.
@@ 4.2 Are the current pixel-level MLLMs heading in the right direction?
In order to answer this, we evaluate each of these pixel-level MLLMs VQA capabilities in challenging tasks. Additionally, we evaluate their ability to visually ground the objects of interest in these questions. Table 1 shows the results on the challenging PixMMVP and PixCV-Bench. From the accuracy of VQA, MLLMs that are not trained with pixel-level grounding surpass their pixel-level counterpart with up to 16%. The best in pixel-level MLLMs score in this aspect is GLAMM-RegCap [27] yet it has degraded ability to generate segmentation. On the other hand, when looking at pixel-level visual grounding we find the best model, GLAMM [16], has a weak ability in VQA or following instructions. Moreover, it shows LISA and Llava-G are mostly incapable of following the instruction to output the option letter reported in, . OMG-Llava strikes the right balance in both VQA and pixel-level grounding with the highest score, , within pixel-level MLLMs. However, looking at the bottom three rows, the oracle, PixFoundation, confirms that MLLMs that were never trained with pixel-level grounding have the correct grounding within their learned attention maps yielding higher score, , than OMG-Llava. Qualitative comparison is provided in Fig. 5 and additional examples are in App. B. Additionally, the automatic baseline outperforms OMG-Llava in PixMMVP with a considerable margin and is on-par to it on PixCV-Bench. Furthermore, the attend and segment baseline [2] lags behind our automatic method with more than 10% in both datasets. Refer to App. C for a failure case analysis.
Finally, we evaluate whether the failures of these MLLMs occur in visual grounding, VQA or both. Figure 6 shows the frequency of failures per category, where the majority stem from failures in both especially in the pixel-level MLLMs. While the vanilla MLLMs perform better in the VQA and tend to fail more in the grounding task.
Summary. In summary, pixel-level grounding supervision degrades MLLMs ability in VQA and sometimes even their generalization in grounding. We show that MLLMs trained with pixel-level supervision lag behind vanilla MLLMs using simple mechanisms to extract grounding, and the oracle indicates there is an opportunity to improve this. Moreover, we show that grounding might not coincide with the noun phrase most similar to the referred expression, where our oracle upper bound and automatic baseline are both surpassing the attend and segment.
@@ 4.3 When does grounding emerge in MLLMs?
When - location. Taking into account the powerful performance of the oracle upper bound, it begs the important question of when grounding emerges. We start by looking at when it emerges in terms of the location within the output text. We analyze the word/phrase location with respect to the full output text in terms of a percentage from its total length, (i.e., 0% means the beginning of the text). Accordingly, Fig. 7(a) shows the location percentages histogram, binned at 10%, for the three base MLLMs reporting the oracle selection and evaluating on PixMMVP benchmark using the second probing. In the Llava 1.5 variants, the highest grounding emerges at the last 40% of the output text, while for Cambrian it emerges at the last 60%.
When - concept. For the second analysis we look into the concept category that the correct output word/phrase corresponds to. The previous assumption in other works is that grounding emerges in the exact noun/noun phrase of the object of interest. Except our analysis confirms that this is not necessarily the case. We take the correct noun/noun phrase where the grounding emerges based on the oracle from all the three variants, then we pass it to GPT-4o to request a grouping of these concepts. It resulted into six main groups, which are: (i) Color and appearance, (ii) location and position, (iii) object parts, (iv) context and setting, (v) objects and entities, and (vi) State. We then prompt for each of the noun/noun phrase, GPT-4o, to categorize it within these six categories. The histogram of the occurrences of these concept categories is shown in Fig. 7(b). It clearly conveys that in certain scenarios the correct output when grounding emerges can be describing the position or the color of the object of interest not necessarily the exact ground-truth referring expression. Fig. 8 shows qualitative examples of these scenarios from PixMMVP using the second probing. Additional results are provided in App. B.
@@
Summary. In summary, we found that highest frequent emergence of grounding is in the last 40-60% of the output text on PixMMVP which might indicate that the reasoning of the model on this challenging task to provide a response impacts when the correct visual grounding can occur. More importantly, we show that grounding in MLLMs can emerge in the noun phrase that corresponds to color, position or other characteristics of the object of interest and not necessarily the exact referring expression.
@@ Appendix A Additional implementation details
Models. We also detail the model checkpoints we use for the four pixel-level MLLMs and their variants, retrieved from HuggingFace [22] in Table 2. These also include the model checkpoints used for the base MLLMs that were not trained with pixel-level visual grounding. It is worth noting, that for GLAMM we use two variants (FullScope and RegCap) since their base model (i.e., FullScope) has low performance in the visual question answering task. As such we use the other variant for GLAMM that was fine-tuned for region-level captioning using RefCOCOg dataset. Furthermore, we provide details on the oracle selection mechanism, we discard the cases where the ground-truth segmentation is all background in the when analysis, since there is no ground-truth grounding to evaluate against. These occur in few cases in PixMMVP.
Additionally, we provide details on the SAM model that is used in the three baselines and upper bounds in our benchmarks, where we use the ViT-H variant. Finally, we provide an illustrative example of our automatic selection mechanism with the corresponding predictions on PixMMVP using Llava 1.5 (7B) in Fig. 11. Our automatic selection goes through an iterative process of prompting the selected MLLM, in our case GPT-4o, with groups of images highlighting the predicted segmentation to select the best within each group. In the final stage the best images are used to prompt the MLLM to select the final selected mask that best describes the object of interest.
@@
| Llava 1.5 (7B) | liuhaotian/llava-v1.5-7b |
| Llava 1.5 (13B) | liuhaotian/llava-v1.5-13b |
@@
Evaluation. We also provide the details on computing the visual question answering accuracy using GPT-4o in the first protocol [19]. We use the following prompt: “Given the following question QUESTION, the correct answer is ANSWER. Does the following answer correctly answers the question, answer: RESPONSE? Respond with a Yes/No”.
@@ B.1 Baselines ablation
Attend and Segment [2] (Cambrian-1 (8B))
@@ B.3 When does grounding emerge?
We show additional examples on when grounding emerges in multi-modal large language models, specifically in the Llava 1.5 (7B) variant, using the second probing to prompt the model to segment what is in the referring expression. Figures 14, 16, 15 and 17 show the corresponding predicted masks for the grounding that emerged highlighted in red with the maximum attention point as a black dot. Figure 8 in the main text shows the aforementioned four examples with the input prompt, the concept category and the noun phrase corresponding to the best grounding using the oracle selection and the full output text. It clearly shows that the correct output token can correspond to location or color but not necessarily the ground-truth referring expression. While some of the noun phrases and their masks from the point prompting of SAM do correspond to what the noun phrase is describing. It is not always the case, example in Fig. 16 “the flame” was not able to highlight the correct object yet it appeared in the noun phrase corresponding to the location “the top”. It also shows that the majority of correct grounding occurs near the end of the output text after the model reasoned for an answer to the question.
Additionally, this type of mining for attention maps can show potential mistakes that MLLMs fall into while performing the grounding to reason on what is in the image. An example on that is in Fig. 17, where the model potentially mistakes a crack in the wood for the minute hand. Thus, our PixFoundation serves as an interesting tool to interpret and understand how MLLMs work and reason to produce the final output with the oracle selection as an upper bound. Yet, there is an open question on how to identify the real mistake from the MLLM vs. simple misalignment between the noun phrase and the grounding emerging, since grounding could emerge with phrases describing it or parts of it instead of the object’s noun phrase itself. We leave this for future work.
@@ B.5 PixCV-Bench benchmark
Appendix C Failure case analysis
C.1 Failures in visual question answering
C.2 Failures in pixel-level visual grounding
Appendix D Additional quantitative analysis
D.1 When grounding emerges - CV-Bench
In Fig. 23(a) we show the analysis on when grounding emerges on PixCV-Bench in terms of the frequency of the grounding location. It is worth noting that PixMMVP is more challenging than PixCV-Bench evidently from the reported IoU and accuracy metrics on both reported in Table 1. It seems on the less challenging dataset PixCV-Bench, grounding tends to emerge frequently near the beginning of the output. This might relate to PixMMVP being more challenging in terms of the level of reasoning than CV-Bench or the fact that PixMMVP poses a harder referring segmentation task than PixCV-Bench, which is mostly using the class names. However, the consistent finding among both datasets is that grounding can emerge coinciding to various concept categories whether location, color or state as shown in Fig. 23(b).
D.2 Analysis on the output length
In this section, we provide additional analysis on the output length on average through PixMMVP dataset using the first and second probing schemes. Specifically, we report the output length as the number of characters in the output, and the number of noun phrases extracted from it. The reason to study this, since it has relation to the number of noun phrases and consequently the number of masks our baselines are selecting among. Table 3 shows the average output length computed across PixMMVP dataset, comparing the three base MLLMs. We notice that Cambrian-1 (8B) generates longer outputs with a considerable margin than Llava variants. Hence, we believe the superiority of the oracle upper bound with Cambrian-1 in the grounding has strong correlation to producing longer outputs with more attention maps to mine and select from, than Llava variants.